OpenStudy (anonymous):

The area of a triangle is 119 square meters. The base is 17 meters long. What is the height of the triangle? A. 3.5 m B. 7 m C. 14 m D. 28 m

OpenStudy (anonymous):

@hybrik

OpenStudy (hybrik):

To find this, plug the formula A = 1/2 BH, 119 = 1/2 17*h

OpenStudy (hybrik):

119 * 2/17 = h simplify

OpenStudy (anonymous):

@yungjvke

OpenStudy (hybrik):

Simplify that to get 14

OpenStudy (anonymous):

thanks